# Haoyu Zhang12  Qiaohui Chu12  Meng Liu3  Yunxiao Wang4  Bin Wen5  Fan Yang5 Tingting Gao5  Di Zhang5  Yaowei Wang12  Liqiang Nie1

### v1

@@ C.1.4 Charades-Ego
Table 15: Performance comparison of Exo2Ego across different training stages on the EgoBench benchmark. The QAEgo4D dataset is categorized into “Closed” and “Open” question answering, while the EgoTaskQA dataset includes “Direct” and “Indirect” subsets. In EgoMCQ, “Inter” refers to cases where the candidates and input clips come from different videos, while “Intra” indicates that they are from the same video, presenting a more challenging scenario. The best results are highlighted in bold.
C.1.5 EPIC-KITCHENS-100
EPIC-KITCHENS-100 [15] is a first-person vision dataset featuring over 100 hours of unscripted kitchen activities recorded in 45 kitchens across 4 cities. It contains 20 million frames, 90,000 action segments, and 20,000 unique activity descriptions, covering 97 verb classes and 300 noun classes. Captured using head-mounted cameras, the dataset supports a wide range of tasks, including action recognition, video classification, and multimodal learning, and is notable for its innovative “Pause-and-Talk” annotation system. This dataset addresses five key challenges in advancing egocentric video understanding.
@@ C.1.5 EPIC-KITCHENS-100
C.1.6 VLN-QA
VLN-QA [38] is a multiple-choice question-answering dataset designed to evaluate the navigation capabilities of vision-language systems in complex indoor environments. Built on the VLN-CE source, it contains thousands of curated question-answer pairs based on first-person video clips that simulate real-world navigation tasks. Each video lasts several minutes and challenges models to make decisions based on multimodal understanding, including visual perception, spatial reasoning, and path planning, all within intricate indoor settings.
This dataset is built by VideoChat2 [43] using LLMs based on VLN-CE. Since it already fulfills the QA format requirements, we directly adopt the processed data provided by VideoChat2.
C.1.7 EgoPlan
EgoPlan [10] is a comprehensive egocentric video benchmark designed to assess human-level planning capabilities in vision and language systems. Built from EPIC-KITCHENS [15] and Ego4D [24], it features 4,939 verified multiple-choice questions, covering 3,296 task goals and 3,185 action plans across 419 real-world scenes. The dataset involves interactions with 558 objects and 234 verbs, challenging models to select the correct next action plan based on visual observations from egocentric video clips. This makes EgoPlan a valuable tool for advancing multimodal models in real-world planning scenarios.
@@ C.1.8 EgoMCQ
EgoMCQ [49] is a multiple-choice question-answering dataset designed to assess video-text alignment in egocentric vision systems. Derived from Ego4D [24], it includes 39,000 questions based on 468 hours of egocentric video covering a wide range of human activities. Each question involves selecting the correct video clip from five options based on a narration, with two settings: “inter-video”, for distinguishing between different videos, and “intra-video”, for fine-grained context within the same video.
@@
This dataset provides complete multiple-choice QA data. We use the video in each query as the visual input, and the original text options as candidates, to construct a video-text retrieval task. Since most methods can achieve 95% accuracy (inter-video) under the original settings (1: 5, a query corresponds to 5 candidates), we increased the number of candidates to 10 (i.e., 1: 10) by adding similar text based on SentenceBERT [69] similarity to highlight the differences among them.
C.2 Example Presentation
@@
In Table 12, we visually present an example from each data source included in our EgoBench benchmark. Each data source emphasizes a different aspect of task evaluation. For instance, QAEgo4D focuses on episodic memory, with questions like “Where did I put the chopsticks?” that require precise recall of past details and comprehension to select the correct answer. On the other hand, the Egoschema source emphasizes the evaluation of the model’s reasoning abilities, requiring a comprehensive understanding of the entire event, including its causes and consequences.
D More Experiments
D.1 Hyperparameter Settings
@@
Table 14 summarizes the hyperparameters used in different training stages. At each stage, we used 16-frame videos with a resolution of 336 $\times$ 336 as the visual input. The overall framework employs the AdamW [37, 94] optimizer with momentum parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$, along with a weight decay of 0.02 to mitigate overfitting. The learning rate follows a cosine decay schedule for gradual reduction. In line with VideoLLaMA2, we enabled BF16 and TF32 precision while disabling FP16 precision to optimize performance using mixed-precision calculations on compatible hardware. To further enhance training efficiency, we implemented flash attention [16].
D.2 Comparison Results
D.2.1 Comparison with Qwen2-VL
Qwen2-VL [81], published in arXiv in October 2024, has obtained optimal performance on multiple existing benchmarks and achieved competitive results with the GPT-4 series. In Table 13, we included its experimental results on the proposed EgoBench and compared them with VideoLLaMA2 [11], VideoChat2 [43], and our Exo2Ego model.
The results show that, with a uniform 16-frame input, Qwen2-VL does not achieve the anticipated outstanding performance on EgoBench, performing similarly to other contemporary baselines and significantly underperforming compared to our Exo2Ego model on most tasks. This result further underscores the differences between egocentric and exocentric data domains and highlights the superiority of our framework for egocentric video understanding.
D.2.2 Effect of stages
@@
To explore the impact of different stages on the overall performance of the model, we evaluated each stage separately and the results are shown in Table 15. From the results, we can observe that pre-training in the initialization provides basic egocentric video comprehension, while mapping learning in the Stage 2 effectively aligns data from different domains, leading to competitive results on EgoBench. After Stage 3, the model’s instruction-following capability is enhanced, enabling it to better understand instructions for downstream tasks, which significantly improves its performance on the EgoBench benchmark. These results fully validate the effect and effectiveness of each stage in the overall pipeline.
D.2.3 Effect of prompts
@@
Figure 14: Additional qualitative examples from the EgoBench benchmark. “GT” denotes the ground truth.
D.3 Qualitative Analysis
We present additional qualitative examples from various data sources in the EgoBench benchmark, as shown in Figures 13 and 14. A detailed analysis is summarized below:
@@ D.3 Qualitative Analysis
E Limitations
The current scale of egocentric data used for training and evaluation is relatively small, with limited diversity, which may lead to insufficient learning and incomplete validation. Additionally, employing more advanced teacher models (e.g., Qwen2-VL [81]) and investigating more effective mapping techniques could facilitate the transfer of more precise exocentric knowledge, thereby enhancing the understanding of egocentric videos.

### v2

@@ Conclusion
Table 9: Performance comparison of Exo2Ego across different training stages on the multiple benchmarks. The QAEgo4D dataset is categorized into “Closed” and “Open” question answering, while the EgoTaskQA dataset includes “Direct” and “Indirect” subsets. In EgoMCQ, “Inter” refers to cases where the candidates and input clips come from different videos, while “Intra” indicates that they are from the same video, presenting a more challenging scenario. The best results are highlighted in bold.
EPIC-KITCHENS-100
EPIC-KITCHENS-100 (damen2022rescaling) is a first-person vision dataset featuring over 100 hours of unscripted kitchen activities recorded in 45 kitchens across 4 cities. It contains 20 million frames, 90,000 action segments, and 20,000 unique activity descriptions, covering 97 verb classes and 300 noun classes. Captured using head-mounted cameras, the dataset supports a wide range of tasks, including action recognition, video classification, and multimodal learning, and is notable for its innovative “Pause-and-Talk” annotation system. This dataset addresses five key challenges in advancing egocentric video understanding.
@@
VLN-QA
VLN-QA (krantz2020beyond) is a multiple-choice question-answering dataset designed to evaluate the navigation capabilities of vision-language systems in complex indoor environments. Built on the VLN-CE source, it contains thousands of curated question-answer pairs based on first-person video clips that simulate real-world navigation tasks. Each video lasts several minutes and challenges models to make decisions based on multimodal understanding, including visual perception, spatial reasoning, and path planning, all within intricate indoor settings.
This dataset is built by VideoChat2 (li2024mvbench) using LLMs based on VLN-CE. Since it already fulfills the QA format requirements, we directly adopt the processed data provided by VideoChat2.
EgoPlan
EgoPlan (chen2023egoplan) is a comprehensive egocentric video benchmark designed to assess human-level planning capabilities in vision and language systems. Built from EPIC-KITCHENS (damen2022rescaling) and Ego4D (grauman2022ego4d), it features 4,939 verified multiple-choice questions, covering 3,296 task goals and 3,185 action plans across 419 real-world scenes. The dataset involves interactions with 558 objects and 234 verbs, challenging models to select the correct next action plan based on visual observations from egocentric video clips. This makes EgoPlan a valuable tool for advancing multimodal models in real-world planning scenarios.
@@
EgoMCQ (lin2022egocentric) is a multiple-choice question-answering dataset designed to assess video-text alignment in egocentric vision systems. Derived from Ego4D (grauman2022ego4d), it includes 39,000 questions based on 468 hours of egocentric video covering a wide range of human activities. Each question involves selecting the correct video clip from five options based on a narration, with two settings: “inter-video”, for distinguishing between different videos, and “intra-video”, for fine-grained context within the same video.
@@
This dataset provides complete multiple-choice QA data. We use the video in each query as the visual input, and the original text options as candidates, to construct a video-text retrieval task. Since most methods can achieve 95% accuracy (inter-video) under the original settings (1: 5, a query corresponds to 5 candidates), we increased the number of candidates to 10 (i.e., 1: 10) by adding similar text based on SentenceBERT (reimers2019sentence) similarity to highlight the differences among them.
Example Presentation
@@
In Table 8, we visually present an example from each benchmark. Each data source emphasizes a different aspect of task evaluation. For instance, QAEgo4D focuses on episodic memory, with questions like “Where did I put the chopsticks?” that require precise recall of past details and comprehension to select the correct answer. On the other hand, the Egoschema source emphasizes the evaluation of the model’s reasoning abilities, requiring a comprehensive understanding of the entire event, including its causes and consequences.
More Experiments
Hyperparameter Settings
@@
Table 11 summarizes the hyperparameters used in different training stages. At each stage, we used 16-frame videos with a resolution of 336 $\times$ 336 as the visual input. The overall framework employs the AdamW (kingma2014adam) optimizer with momentum parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$, along with a weight decay of 0.02 to mitigate overfitting. The learning rate follows a cosine decay schedule for gradual reduction. In line with VideoLLaMA2, we enabled BF16 and TF32 precision while disabling FP16 precision to optimize performance using mixed-precision calculations on compatible hardware. To further enhance training efficiency, we implemented flash attention (dao2022flashattention).
Comparison Results
Effect of stages
@@
To explore the impact of different stages on the overall performance of the model, we evaluated each stage separately and the results are shown in Table 9. From the results, we can observe that pre-training in the initialization provides basic egocentric video comprehension, while mapping learning in the Stage 2 effectively aligns data from different domains, leading to competitive results on all benchmarks. After Stage 3, the model’s instruction-following capability is enhanced, enabling it to better understand instructions for downstream tasks, which significantly improves its performance. These results fully validate the effect and effectiveness of each stage in the overall pipeline.
Effect of prompts
@@
Figure 14: Additional qualitative examples from all benchmarks. “GT” denotes the ground truth.
Config Initial Stage 1&2 Stage 3
Global Batch size 512 256 64
Local Batch size 8 8 8
Learning rate 1e-3 1e-4 2e-5
Warmup ratio 0.1 0.03 0.03
Total epoch 5 2 1
Table 11: Hyperparameter settings for different training stages.
Qualitative Analysis
We present additional qualitative examples from various data sources, as shown in Figures 13 and 14. A detailed analysis is summarized below:
@@
Limitations
The current scale of egocentric data used for training and evaluation is relatively small, with limited diversity, which may lead to insufficient learning and incomplete validation. Additionally, employing more advanced teacher models (e.g., Qwen2-VL (wang2024qwen2)) and investigating more effective mapping techniques could facilitate the transfer of more precise exocentric knowledge, thereby enhancing the understanding of egocentric videos.